OpenStudy (anonymous):

Find the production level that minimizes the average cost. c(x) = 10e^0.02x

OpenStudy (anonymous):

@amistre64 can u help?

OpenStudy (amistre64):

the word average implies that we divide by c(x) by x

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

so we would want the deriavtive of C(x)/x:\[\frac{xC'-x'C}{x^2}=0\]

OpenStudy (anonymous):

okay so 10e^002x -10.202 / x^2 ?

OpenStudy (amistre64):

something like that yes, but the x^2 is useless to us since \(\frac 0n=0\) so lets focus on the top \[xC'-C=0\] \[ .2x e^{0.02x}-10e^{0.02x}=0\] since e^whatever is never 0, lets factor it out and solve\[.2x-10=0\]

OpenStudy (amistre64):

its either that, or when the derivaitve is undefined ... namely at x=0, but that is not a good value to use in any case

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x = 50 i got after i solved

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

okay